# IQ-Aware Precoding for Atomic MIMO Receivers

### v1

@@ II Review of Rydberg Atomic MIMO System
III Mutual Information of Atomic MIMO Systems
In this section, we introduce a strong reference approximation to linearize the non-linear transition model in (5), built on which the mutual information, $\mathcal{I}({\mathbf{y}};{\mathbf{x}})$, is explicitly formulated. We then analyze the DoF of atomic MIMO systems relying on the modeled mutual information.
III-A Approximated Mutual Information
It is evident from (5) that the atomic receiver exhibits a non-linear input-output relationship, given that Rabi frequencies, $\{\Omega_{m}\}$, only capture the magnitude of incident EM waves. This fact makes it intractable to express the mutual information, $\mathcal{I}({\mathbf{y}};{\mathbf{x}})$, analytically, let alone maximizing it. To overcome this challenge, we note that the artificially introduced LOs are typically placed close to the atomic receiver in practical implementations [17, 18, 19, 32]. In particular, the LO-to-receiver distance is almost tens of centimeters, which is significantly shorter than the transmitter-to-receiver distance. Considering the quadratic-scaling path loss [33], the reference signal can be much stronger than the wireless data and noise. This strong reference approximation allows us to linearize the transmission model in (5) as follows.
Proposition 1.
(Strong reference approximation): If each entry of the reference signal is greatly stronger than the wireless data and noise, i.e.,
$|r_{m}|\gg|{\mathbf{H}}_{m,:}{\mathbf{x}}+w_{m}|,$ (7)
then we can approximate ${\mathbf{y}}$ as
$\boxed{\tilde{{\mathbf{y}}}={\mathbf{y}}-|{\mathbf{r}}|\approx\mathrm{Re}(\tilde{{\mathbf{H}}}{\mathbf{x}})+{\mathbf{w}}_{I},}$ (8)
where ${\mathbf{w}}_{I}\sim\mathcal{N}(0,\frac{\sigma^{2}}{2}{\mathbf{I}}_{N_{r}})$ and $\tilde{{\mathbf{H}}}_{m,:}\overset{\Delta}{=}e^{-j\angle r_{m}}{\mathbf{H}}_{m,:}$.
@@ III-A Approximated Mutual Information
Proposition 1 indicates that, in the strong-reference-signal regime, the non-linear magnitude detector in (5) can be transformed into a linear real-part detector in (8) by subtracting the offset $|{\mathbf{r}}|$. More importantly, the complicated mutual information, $\mathcal{I}({\mathbf{y}};{\mathbf{x}})=\mathcal{I}(\tilde{{\mathbf{y}}};{\mathbf{x}})$, is converted to a tractable expression:
$\mathcal{I}(\tilde{{\mathbf{y}}};{\mathbf{x}})\rightarrow\mathcal{I}(\mathrm{Re}(\tilde{{\mathbf{H}}}{\mathbf{x}})+{\mathbf{w}}_{I};{\mathbf{x}}).$ (9)
To further simplify (8) and (9), we introduce the subscript $I$ and $Q$ to denote the real (in-phase) and imaginary (quadrature) components of a complex number, e.g., $\tilde{{\mathbf{H}}}=\tilde{{\mathbf{H}}}_{I}+j\tilde{{\mathbf{H}}}_{Q}$ and ${\mathbf{x}}={\mathbf{x}}_{I}+j{\mathbf{x}}_{Q}$. Then, the received signal $\tilde{{\mathbf{y}}}$ is rewritten as
$\tilde{{\mathbf{y}}}=\underbrace{(\tilde{{\mathbf{H}}}_{I},-\tilde{{\mathbf{H}}}_{Q})}_{\bar{{\mathbf{H}}}}\underbrace{\left(\begin{array}[]{c}{\mathbf{x}}_{I}\\ {\mathbf{x}}_{Q}\end{array}\right)}_{\bar{{\mathbf{x}}}}+{\mathbf{w}}_{I},$ (12)
where $\bar{{\mathbf{H}}}\in\mathbb{R}^{N_{r}\times 2N_{t}}$ and $\bar{{\mathbf{x}}}\in\mathbb{R}^{2N_{t}\times 1}$ are defined as the equivalent real-valued channel and radiated signal, respectively. Attributed to the linear behavior of real-part detectors, the maximal throughput is achieved when the radiated signal $\bar{{\mathbf{x}}}$ follows the Gaussian distribution $\bar{{\mathbf{x}}}\sim\mathcal{N}(0,\bar{{\mathbf{Q}}})$, with the covariance matrix $\bar{{\mathbf{Q}}}=\mathsf{E}(\bar{{\mathbf{x}}}\bar{{\mathbf{x}}}^{T})\in\mathbb{R}^{2N_{t}\times 2N_{t}}$. Thereafter, the mutual information is expressed as
$\mathcal{I}(\mathrm{Re}(\tilde{{\mathbf{H}}}{\mathbf{x}})+{\mathbf{w}}_{I};{\mathbf{x}})=\frac{1}{2}\log\det({\mathbf{I}}_{N_{r}}+\frac{2}{\sigma^{2}}\bar{{\mathbf{H}}}\bar{{\mathbf{Q}}}\bar{{\mathbf{H}}}^{T}).$ (13)
Here, the discount factor $\frac{1}{2}$ arises because the variables $\bar{{\mathbf{H}}}$, $\bar{{\mathbf{x}}}$, and ${\mathbf{w}}_{I}$ are all real valued. Moreover, considering the total power constraint imposed on the radiated signal, we have
@@
where $P$ is the transmission power. In summary, leveraging the strong reference approximation, the achievable capacity, $C$, in (6) can be reformulated as
@@
$C=\max_{\mathsf{Tr}\{\bar{{\mathbf{Q}}}\}=P,\>\bar{{\mathbf{Q}}}\succeq 0}\frac{1}{2}\log\det\left({\mathbf{I}}_{N_{r}}+\frac{2}{\sigma^{2}}\bar{{\mathbf{H}}}\bar{{\mathbf{Q}}}\bar{{\mathbf{H}}}^{T}\right).$ (15)
The major difference between the mutual information (15) and that of the conventional MIMO [28] is due to the conversion of channel matrix and covariance matrix from their complex form to real form, i.e., ${\mathbf{H}}\in\mathbb{C}^{N_{r}\times N_{t}}\Rightarrow\bar{{\mathbf{H}}}\in\mathbb{C}^{N_{r}\times 2N_{t}}$ and ${\mathbf{Q}}=\mathsf{E}({\mathbf{x}}{\mathbf{x}}^{H})\Rightarrow\bar{{\mathbf{Q}}}=\mathsf{E}(\bar{{\mathbf{x}}}\bar{{\mathbf{x}}}^{T})$. Therefore, the transmission model (8) prefers real-valued matrix manipulation, rather than complex-valued one.
III-B Degree-of-Freedom Analysis
It is well-known that, at high SNR, the ergodic capacity of a conventional MIMO system is $\min\{N_{r},N_{t}\}\log\frac{\rm SNR}{N_{t}}+\mathcal{O}(1)$, where $\min\{N_{r},N_{t}\}$ signifies the system’s DoF, and the average SNR is defined as ${\rm SNR}=\frac{P}{\sigma^{2}}$ (see [28, 29]). Following the same analysis method, we will study the ergodic capacity of atomic MIMO receivers and the corresponding DoF based on the transition model (8).
The channel state information (CSI) is assumed to be known to the receiver, but not at the transmitter [29]. Given this assumption, the covariance matrix of radiated signal should have a uniform power allocation, i.e., $\bar{{\mathbf{Q}}}=\frac{P}{2N_{t}}{\mathbf{I}}_{2N_{t}}$. We assume that the coefficients of ${\mathbf{H}}$ follow an independent and identical distributed (i.i.d) complex circularly symmetric Gaussian distribution with unit variance. The ergodic capacity of this transmission model is then obtained by averaging ${\mathbf{H}}$ (or $\bar{{\mathbf{H}}}$):
$C_{e}=\mathsf{E}\left[\frac{1}{2}\log\det\left({\mathbf{I}}_{N_{r}}+\frac{\rm SNR}{N_{t}}\bar{{\mathbf{H}}}\bar{{\mathbf{H}}}^{T}\right)\right].$ (16)
Following the standard analysis of MIMO channel capacity, the asymptotic behavior of $C_{e}$ at high SNRs is derived below.
Proposition 2.
At high SNRs, the ergodic capacity, $C_{e}$, is asymptotically
$\boxed{C_{e}=\min\left\{\frac{N_{r}}{2},N_{t}\right\}\log\frac{\mathrm{SNR}}{N_{t}}+\mathcal{O}(1),\quad(\text{SNR}\rightarrow\infty).}$ (17)
Proof: (See Appendix B).  $\Box$
Figure 2: Comparison between the three DoFs: $\frac{1}{2}\min\{N_{r},N_{t}\}$, $\min\{\frac{N_{r}}{2},N_{t}\}$, and $\min\{N_{r},N_{t}\}$, where $N_{t}=5$.
Proposition 2 suggests that the transmission rate of atomic MIMO system is improved by $\min\{N_{r}/2,N_{t}\}$ bit/s/Hz for every 3 dB increment in SNR, which corresponds to a DoF of $\lim_{\text{SNR}\rightarrow\infty}\frac{C_{e}}{\log\text{SNR}}=\min\left\{\frac{N_{r}}{2},N_{t}\right\}$. This result is attributed to the fact that the DoF is lost at the receiver side due to the absence of imaginary-part information, while the DoF remains intact at the transmitter side. Thereafter, a discount factor $\frac{1}{2}$ is imposed on $N_{r}$, giving rise to a DoF of $\min\{\frac{N_{r}}{2},N_{t}\}$. To clarify our discussion, we compare three DoFs, $\min\{N_{r},N_{t}\}$, $\min\{\frac{N_{r}}{2},N_{t}\}$, and $\frac{1}{2}\min\{N_{r},N_{t}\}$, as a function of $N_{r}$ in Fig. 2. These three DoFs have their own physical meaning: the former two refer to DoFs of conventional MIMO and atomic MIMO systems, while the last one is interpreted as the DoF of real-valued symbols transmission on the in-phase channel.
As $N_{r}$ increases, the DoF $\min\{\frac{N_{r}}{2},N_{t}\}$ gradually climbs from $\frac{1}{2}\min\{N_{r},N_{t}\}$ to $\min\{N_{r},N_{t}\}$. Specifically, when $N_{r}<N_{t}$, we have $\min\left\{\frac{N_{r}}{2},N_{t}\right\}=\frac{1}{2}\min\left\{{N_{r}},N_{t}\right\}=\frac{1}{2}N_{r}$. The DoF of atomic MIMO systems is exactly half of that of conventional systems. In this context, it is sufficient to achieve the capacity of atomic receivers by transmitting $\frac{N_{r}}{2}$ complex-valued symbols, or equivalently, $N_{r}$ real-valued symbols on the in-phase channel only. Moreover, as $N_{r}$ grows from $N_{t}$ to $2N_{t}$, the DoF, $\min\{N_{r},N_{t}\}$, get saturated while the gap between $\min\{N_{r},N_{t}\}$ and $\min\{N_{r}/2,N_{t}\}$ narrows. Particularly, when $N_{r}\geq 2N_{t}$, we have $\min\{\frac{N_{r}}{2},N_{t}\}=\min\{N_{r},N_{t}\}=N_{t}$. This implies that an atomic receiver does not suffer any DoF loss compared to its traditional counterpart if the number of atomic antennas is at least twice the number of transmit antennas. This is because the $N_{r}$ real-valued observations are sufficient to recover $N_{t}\leq\frac{N_{r}}{2}$ complex-valued symbols. Consequently, under the same high SNR condition, the capacity gap between atomic MIMO and classic MIMO systems tends to a constant independent of SNR. In summary, to fully unleash the multiplexing gain of an atomic MIMO system, it is necessary to have $N_{r}\geq 2N_{t}$.
@@ IV IQ-Aware Digital Precoding
When channel state information is available at both receiver and transmitter ends, we can manipulate the covariance matrix $\bar{{\mathbf{Q}}}$ on purpose using precoding technique to achieve the instantaneous capacity in (15) for each channel realization. In this section, we propose an IQ-aware precoding approach designed to optimize digital precoding for atomic MIMO receivers. A comparison with traditional digital precoding is also presented for a comprehensive understanding.
IV-A Precoding architecture
Figure 3: (a) The classic complex digital precoding and (b) the proposed IQ-aware digital precoding.
We define ${\mathbf{s}}_{I}\in\mathbb{R}^{N_{s}}$ and ${\mathbf{s}}_{Q}\in\mathbb{R}^{N_{s}}$ as the transmitted symbols on the in-phase and quadrature (IQ) channels, respectively. The variable $N_{s}$ represents the number of multiplexed data streams, which should be no larger than the number of maximum parallel spatial channels, i.e., $N_{s}\leq\min\{N_{r}/2,N_{t}\}$. The overall baseband symbol is denoted as $\bar{{\mathbf{s}}}=[{\mathbf{s}}_{I}^{T},{\mathbf{s}}_{Q}^{T}]^{T}\in\mathbb{R}^{2N_{s}}$ in the real form and ${\mathbf{s}}={\mathbf{s}}_{I}+j{\mathbf{s}}_{Q}\in\mathbb{C}^{N_{s}}$ in the complex form. The symbols ${\mathbf{s}}_{I}$ and ${\mathbf{s}}_{Q}$ follow an i.i.d Gaussian distribution $\mathcal{N}(0,\frac{1}{2}{\mathbf{I}}_{N_{s}})$. The objective of precoding is to design a linear mapping $f:\bar{{\mathbf{s}}}\mapsto\bar{{\mathbf{x}}}$ such that the covariance matrix $\bar{{\mathbf{Q}}}$ can maximize the mutual information $\mathcal{I}({\mathbf{y}};{\mathbf{x}})$ as presented in (15).
To this end, we start with briefly reviewing the implementation of classical digital precoding. Classic MIMO systems widely adopt a complex-valued digital precoder ${\mathbf{F}}={\mathbf{F}}_{I}+j{\mathbf{F}}_{Q}\in\mathbb{C}^{N_{t}\times N_{s}}$ as depicted in Fig. 3 (a). The real-valued symbols ${\mathbf{s}}_{I}$ and ${\mathbf{s}}_{Q}$ are jointly processed by the two matrices ${\mathbf{F}}_{I}$ and ${\mathbf{F}}_{Q}$ on the digital baseband before entering the IQ modulator. After combining the IQ components, the transmitted signal in the complex form is thereby the complex product of ${\mathbf{F}}$ and ${\mathbf{s}}$, specifically
${\mathbf{x}}={\mathbf{F}}{\mathbf{s}}={\mathbf{F}}_{I}{\mathbf{s}}_{I}-{\mathbf{F}}_{Q}{\mathbf{s}}_{Q}+j({\mathbf{F}}_{Q}{\mathbf{s}}_{I}+{\mathbf{F}}_{I}{\mathbf{s}}_{Q}).$ (18)
Because classical RF antennas can access both the real and imaginary parts of received signal, the complex-valued precoding is able to achieve the instantaneous capacity of classical MIMO by assigning ${\mathbf{F}}$ with the principal singular vectors of the complex channel, ${\mathbf{H}}$. However, this method is infeasible for atomic MIMO systems. Specifically, to maximize the mutual information (15), it is necessary to expand the complex signal ${\mathbf{x}}$ to the real form:

### v2

@@ II Review of Rydberg Atomic MIMO System
III Channel Capacity With Atomic MIMO Receivers
In this section, we introduce a strong-reference approximation to linearize the non-linear transition model in (5). This allows the mutual information, $\mathcal{I}({\mathbf{y}};{\mathbf{x}})$, as well as the channel capacity, $C$, to be explicitly formulated.
For conventional MIMO systems, both the amplitude and phase information of received signal are accessible, resulting in a linear transition model: ${\mathbf{y}}={\mathbf{H}}{\mathbf{x}}+{\mathbf{w}}$. It has been proven that the channel capacity, $C^{\prime}$, of this linear MIMO model is achievable if and only if ${\mathbf{x}}$ follows a circularly-symmetric complex Gaussian (CSCG) distribution [^1] A Gaussian random vector ${\mathbf{x}}$ is CSCG if and only if $e^{j\theta}{\mathbf{x}}$ has the same distribution with ${\mathbf{x}}$ for any real scalar $\theta$., i.e.,
$C^{\prime}=\max_{{\mathbf{x}}\sim\mathcal{CN}(0,{\mathbf{Q}})}\log\det\left({\mathbf{I}}_{N_{r}}+\frac{1}{\sigma^{2}}{\mathbf{H}}{\mathbf{Q}}{\mathbf{H}}^{H}.\right),$ (7)
where ${\mathbf{Q}}=\mathsf{E}({\mathbf{x}}{\mathbf{x}}^{H})$. For an atomic MIMO receiver, however, it is evident from (5) that it exhibits a non-linear input-output relationship, given that Rabi frequencies, $\{\Omega_{m}\}$, only capture the magnitude of incident EM waves. This fact makes it intractable to express the mutual information, $\mathcal{I}({\mathbf{y}};{\mathbf{x}})$, analytically, let alone calculating the channel capacity, $C$.
To overcome the above challenge, we notice that the LO (see Fig. 1) is typically placed close to the atomic receiver [18, 19, 20, 34]. As a result, the LO-to-receiver distance (e.g., tens of centimeters) is much shorter than the transmitter-to-receiver distance (e.g., tens of meters), resulting in a much stronger reference signal than the wireless signal and noise. This admits a strong-reference approximation to linearize the transmission model in (5) as follows.
Lemma 1.
(Strong-reference approximation): If each entry of the reference signal is greatly stronger than the wireless signal and noise, i.e., $|r_{m}|\gg|{\mathbf{H}}_{m,:}{\mathbf{x}}+w_{m}|$, we can approximate ${\mathbf{y}}$ in (5) as
$\boxed{{\mathbf{y}}-|{\mathbf{r}}|\approx\bar{{\mathbf{y}}}=\mathrm{Re}(\widetilde{{\mathbf{H}}}{\mathbf{x}})+\bar{{\mathbf{w}}},}$ (8)
where $\widetilde{{\mathbf{H}}}_{m,:}\overset{\Delta}{=}e^{-j\angle r_{m}}{\mathbf{H}}_{m,:}$ and $\bar{{\mathbf{w}}}\sim\mathcal{N}(0,\frac{\sigma^{2}}{2}{\mathbf{I}}_{N_{r}})$.
@@ III Channel Capacity With Atomic MIMO Receivers
Lemma 1 indicates that, in the strong-reference-signal regime, the non-linear magnitude detector in (5) can be transformed into a linear real-part detector in (8) by subtracting the offset $|{\mathbf{r}}|$. More importantly, the complicated mutual information, $\mathcal{I}({\mathbf{y}};{\mathbf{x}})=\mathcal{I}({{\mathbf{y}}}-|{\mathbf{r}}|;{\mathbf{x}})$, is converted to a tractable expression: $\mathcal{I}({{\mathbf{y}}}-|{\mathbf{r}}|;{\mathbf{x}})\rightarrow\mathcal{I}(\bar{{\mathbf{y}}};{\mathbf{x}})$. To write $\mathcal{I}(\bar{{\mathbf{y}}};{\mathbf{x}})$ analytically, we introduce the subscripts $I$ and $Q$ to denote the real (in-phase) and imaginary (quadrature) components of a complex number, e.g., $\widetilde{{\mathbf{H}}}=\widetilde{{\mathbf{H}}}_{I}+j\widetilde{{\mathbf{H}}}_{Q}$ and ${\mathbf{x}}={\mathbf{x}}_{I}+j{\mathbf{x}}_{Q}$. Then, the received signal $\bar{{\mathbf{y}}}$ in (8) is rewritten as
$\bar{{\mathbf{y}}}=(\widetilde{{\mathbf{H}}}_{I},-\widetilde{{\mathbf{H}}}_{Q})\left(\begin{array}[]{c}{\mathbf{x}}_{I}\\ {\mathbf{x}}_{Q}\end{array}\right)+{\mathbf{w}}_{I}={\bar{{\mathbf{H}}}}{\bar{{\mathbf{x}}}}+\bar{{\mathbf{w}}},$ (11)
where $\bar{{\mathbf{H}}}\overset{\Delta}{=}(\widetilde{{\mathbf{H}}}_{I},-\widetilde{{\mathbf{H}}}_{Q})\in\mathbb{R}^{N_{r}\times 2N_{t}}$ and $\bar{{\mathbf{x}}}\overset{\Delta}{=}({{\mathbf{x}}}_{I}^{T},{{\mathbf{x}}}_{Q}^{T})^{T}\in\mathbb{R}^{2N_{t}\times 1}$ represent the real-valued channel and radiated signal, respectively. Attributed to the linear behavior of real-part detectors, the maximal throughput is achieved when the radiated signal $\bar{{\mathbf{x}}}$ follows the Gaussian distribution $\bar{{\mathbf{x}}}\sim\mathcal{N}(0,\bar{{\mathbf{Q}}})$, with the covariance matrix $\bar{{\mathbf{Q}}}=\mathsf{E}(\bar{{\mathbf{x}}}\bar{{\mathbf{x}}}^{T})\in\mathbb{R}^{2N_{t}\times 2N_{t}}$. Thereafter, the mutual information is expressed as
$\mathcal{I}(\bar{{\mathbf{y}}};{\mathbf{x}})=\frac{1}{2}\log\det({\mathbf{I}}_{N_{r}}+\frac{2}{\sigma^{2}}\bar{{\mathbf{H}}}\bar{{\mathbf{Q}}}\bar{{\mathbf{H}}}^{T}).$ (12)
Here, the discount factor $\frac{1}{2}$ arises because the variables $\bar{{\mathbf{H}}}$, $\bar{{\mathbf{x}}}$, and $\bar{{\mathbf{w}}}$ are all real valued. Moreover, considering the total power constraint imposed on the radiated signal, we have
@@
where $P$ is the transmission power. In summary, the channel capacity, $C$, of atomic MIMO systems can be formulated as
@@
$\boxed{C=\max_{\mathsf{Tr}\{\bar{{\mathbf{Q}}}\}=P,\>\bar{{\mathbf{Q}}}\succeq 0}\frac{1}{2}\log\det\left({\mathbf{I}}_{N_{r}}+\frac{2}{\sigma^{2}}\bar{{\mathbf{H}}}\bar{{\mathbf{Q}}}\bar{{\mathbf{H}}}^{T}\right).}$ (14)
Remark 1.
By comparing the channel capacities of conventional MIMO in (7) and atomic MIMO in (14), one can observe that the channel matrix is converted from a complex form, ${\mathbf{H}}$, to a real form, $\bar{{\mathbf{H}}}$. Moreover, the latter capacity is directly related to the radiated signals ${\mathbf{x}}_{I}$ and ${\mathbf{x}}_{Q}$ via their joint distribution, $\bar{{\mathbf{x}}}\sim\mathcal{N}(0,\bar{{\mathbf{Q}}})$, rather than their complex form, ${\mathbf{x}}\sim\mathcal{CN}(0,{{\mathbf{Q}}})$. These observations suggest that achieving the channel capacity of atomic MIMO requires that ${\mathbf{x}}_{I}$ and ${\mathbf{x}}_{Q}$ (or equivalently ${\mathbf{x}}={\mathbf{x}}_{I}+j{\mathbf{x}}_{Q}$) follow a general Gaussian distribution, instead of a CSCG distribution in (7). This difference leads to a distinguishing feature of IQ-aware precoding for atomic MIMO.
@@ IV IQ-Aware Digital Precoding
In this section, we propose an IQ-aware digital precoding method to achieve the channel capacity, $C$, in (14) of atomic MIMO systems. The asymptotic channel capacity at high SNRs is subsequently analyzed to unveil the system DoF.
IV-A Precoding Design
We define ${\mathbf{s}}_{I}\in\mathbb{R}^{N_{s}}$ and ${\mathbf{s}}_{Q}\in\mathbb{R}^{N_{s}}$ as the data symbols in the in-phase and quadrature channels, respectively, where $N_{s}$ is the number of complex data streams. The overall baseband data is denoted as $\bar{{\mathbf{s}}}=[{\mathbf{s}}_{I}^{T},{\mathbf{s}}_{Q}^{T}]^{T}\in\mathbb{R}^{2N_{s}}$ in the real form and ${\mathbf{s}}={\mathbf{s}}_{I}+j{\mathbf{s}}_{Q}\in\mathbb{C}^{N_{s}}$ in the complex form. The symbols ${\mathbf{s}}_{I}$ and ${\mathbf{s}}_{Q}$ follow i.i.d Gaussian distributions $\mathcal{N}(0,\frac{1}{2}{\mathbf{I}}_{N_{s}})$, resulting in the CSCG distribution of ${\mathbf{s}}$. The objective of precoding is to design a linear mapping $f:\bar{{\mathbf{s}}}\mapsto\bar{{\mathbf{x}}}$ such that the covariance matrix $\bar{{\mathbf{Q}}}$ can maximize the mutual information in (14).
IV-A1 Sub-optimality of traditional digital precoding
To achieve the channel capacity of a MIMO system, the most intuitive method is to implement traditional complex-valued digital precoding as it has been demonstrated to be optimal in traditional MIMO systems. This, however, is sub-optimal in the case of an atomic MIMO system. Specifically, the classical precoding uses a complex matrix, ${\mathbf{F}}={\mathbf{F}}_{I}+j{\mathbf{F}}_{Q}\in\mathbb{C}^{N_{t}\times N_{s}}$, to jointly process the baseband symbols ${\mathbf{s}}_{I}$ and ${\mathbf{s}}_{Q}$ before the IQ modulator, as shown in Fig. 2 (a). After merging the IQ components, the transmit signal in the complex form is thereby the product of ${\mathbf{F}}$ and ${\mathbf{s}}$:
${\mathbf{x}}={\mathbf{F}}{\mathbf{s}}={\mathbf{F}}_{I}{\mathbf{s}}_{I}-{\mathbf{F}}_{Q}{\mathbf{s}}_{Q}+j({\mathbf{F}}_{Q}{\mathbf{s}}_{I}+{\mathbf{F}}_{I}{\mathbf{s}}_{Q}),$ (15)
whose covariance is ${\mathbf{Q}}=\mathsf{E}({\mathbf{x}}{\mathbf{x}}^{H})={\mathbf{F}}{\mathbf{F}}^{H}$. For subsequent exposition, it is useful to write ${\mathbf{x}}$ in (15) as the equivalent real form: